OpenStudy (anonymous):

Find S22 for -6 + (-4) + (-2) + 0 …

OpenStudy (anonymous):

@summer1nthec1ty did u understand by seeing that link

OpenStudy (anonymous):

no that did nothing...

OpenStudy (anonymous):

then @Jonask will explain .i think he is good at explanation

OpenStudy (anonymous):

the first term is a=-6,and the difference is -4-(-6)=2 the formular for any arithmetic sequence is

OpenStudy (anonymous):

Do u knw to find sum of an ap

OpenStudy (anonymous):

@summer1nthec1ty

OpenStudy (anonymous):

yahoo stop you're embarrassing yourself

OpenStudy (anonymous):

in that link u can click show steps so that u can clear ur doubts

OpenStudy (anonymous):

d = t2 -t1

OpenStudy (anonymous):

\[ S _{n}=n/2[2a+d(n-1)]\]

OpenStudy (anonymous):

JONASK i hate to say it but these equations are not helping me at all...

OpenStudy (anonymous):

wat do u mean @summer1nthec1ty do u want me to go

OpenStudy (anonymous):

do you know the answer, if not then please go (:

OpenStudy (anonymous):

a. 330 b. 332 c. 334 d. 336 ^^ those are the answers which is the right one

OpenStudy (anonymous):

@Jonask plzz dont spell the answer make her to do

OpenStudy (anonymous):

\[S _{22}=22/2[2(-6)+(22-1)(2)]=11(30)\] use your calculator to substitute

OpenStudy (ash2326):

@summer1nthec1ty here on Open Study, we just don't give answers. Please try to understand the solution. Just answers won't help you much. Thanks